OpenStudy (anonymous):

circular cone help!!! radius is 2 L.A- 12 whats the volume and the height????

OpenStudy (anonymous):

if i knew i would help you. you r amazig!!

OpenStudy (anonymous):

lol thank you lol

OpenStudy (anonymous):

What does L.A mean?

OpenStudy (anonymous):

lateral area dear:)

OpenStudy (anonymous):

 apply pythagoras theorem to find height then the volume= 1/3 * base area*height

OpenStudy (shubhamsrg):

LA=12=2(pi)(r)(l) => from here you get l now l= ((h^2)+(r^2)))^(1/2) => from here you get h now volume = (pi)(r^2)(h)/3 hope this helps..

OpenStudy (anonymous):

lateral area? lol, then i am wrong..

OpenStudy (anonymous):

i got h= 4\[\sqrt{8}\]

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

You won't have a rational number as your answer.

OpenStudy (anonymous):

@shubhamsrg , ur formula of latent area is wrong

OpenStudy (shubhamsrg):

oh am sorry,,"2" wont be there..its pi(r)(l)

OpenStudy (anonymous):

since LA=pi.r.slant.h=12 since r=2 so: slant.h=12/2pi=1.9 so the TA=4.pi.r(slant.h+r)=8pi(1.9+2)=98.01 so the Volume=T.A . H but i can't find H

OpenStudy (anonymous):

lateral area = pi*(diameter)*(slant height)